OpenStudy (anonymous):

solve the system by the linear combination method a - 2b = 2 a + b = 2 (b+3)

OpenStudy (anonymous):

which 1 am r u trying to figure out

OpenStudy (anonymous):

both

OpenStudy (whpalmer4):

Multiply the second equation by -1, and add it to the first. This will give you an equation that only contains b. Solve for b, then put the value for b in either of your original equations and solve for a.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry im confused

OpenStudy (whpalmer4):

Equation 1: \[a-2b = 2\]Equation 2:\[a+b = 2(b+3)\] Notice that the coefficient of a in both equations is just 1, so if we take the second equation and subtract it from the first, the a's will drop out and we'll be left with nothing but b and numbers. \[a - 2b - a - b = 2 - 2(b+3)\](I've written the first equation and subtracted out the corresponding sides of the second equation, in case what I did isn't clear) \[-3b = 2 - 2b - 6 = -4b -4\]\[-3b + 4b = -4b -4 + 4b\]\[b=-4\] Now use b = -4 in one of the original equations to solve for a.